OpenStudy (anonymous):

find the indefinite integral (3x-7)^-6 dx

sam (.sam.):

What do you think?

OpenStudy (anonymous):

That I have no idea how to even start! Everything I try I get the wrong answer

sam (.sam.):

Okay we can use power rule for integration here directly, \[\int\limits x^n dx=\frac{x^{n+1}}{n+1}\]

sam (.sam.):

Followed by chain rule \[\int\limits (3x-7)^{-6} dx \\ \\ =\frac{(3x-7)^{-5}}{-5(3)}+c\] The 3 from the denominator comes from chain rule, which means you'll differentiate (3x-7)